Two angles are compementary. The sum of the measure of the first angle and one-fourth the second angle is 66degrees. Find the measures of the angles. What is the measure of the smaller angle? What is the measure of the other angle?
Answer by stanbon(75887)

Two angles are complementary.
x and 90-x
------------------
The sum of the measure of the first angle and one-fourth the second angle is 66 degrees.
90-x + (1/4)x = 66
(-3/4)x = -24
x = 32 degrees
90-x = 58 degrees
----------------------
Find the measures of the angles.
What is the measure of the smaller angle?
What is the measure of the other angle?
-------
Cheers,
Stan H.